# Quantum Circuit Optimization by Graph Coloring

### v1

@@ 4.2 Depth-qubit tradeoffs for QFT-based addition
Figure 3: Depth-qubit tradeoffs for QFT-based addition circuits (a) $\lvert a\rangle\lvert b\rangle\mapsto\lvert a\rangle\lvert b+a\rangle$ and (b) $\lvert a\rangle\lvert b\rangle\lvert c\rangle\mapsto\lvert a\rangle\lvert b\rangle\lvert c+ab\rangle$. The circled and squared points are the most efficient circuit designs with respect to the cost metrics in Eq. (2), respectively (see, Table 1).
@@
An algorithm for the parallelization by extra qubits can be devised by using graph coloring. A heuristic algorithm is developed in which the central idea is noted below. (We do not elaborate on the algorithm as the algorithm itself is not the key founding.)
@@
Consider a commuting circuit and its corresponding graph $G=(V,E)$. Since gates are vertices, introducing extra qubits leaves $V$ unchanged, but depending on which gates are lifted, $E$ changes such that the number of connected edges decreases. Figure 5 shows how the graph can be altered by one extra qubit. Let $G(n,p)$ be a random graph, where $n$ is the number of vertices and $p$ is the probability that an edge exists between two vertices. Bollobás has shown that the chromatic number of a random graph $\chi(G(n,p))$ asymptotically reads [35],
$\chi(G(n,p))\sim\frac{n}{2\log_{1/(1-p)}n}(1+o(1))\enspace.$
In other words, the chromatic number of a random graph gets smaller as $p$ approaches 0. Therefore, the key idea of the algorithm is that the extra qubits are used in a way that the altered graph becomes as sparse in edge density as possible. By using the algorithm, we have parallelized the QFT-based additions $\lvert a\rangle\lvert b\rangle\mapsto\lvert a\rangle\lvert b+a\rangle$ and $\lvert a\rangle\lvert b\rangle\lvert c\rangle\mapsto\lvert a\rangle\lvert b\rangle\lvert c+ab\rangle$. For simplicity, we set $\lceil\log_{2}a\rceil=\lceil\log_{2}b\rceil=n$ and $c=0$. The results are summarized in Figure 3.
@@
$S(C)=\mathsf{depth}(C)\cdot\mathsf{qubit}(C),~{}~{}~{}~{}T(C)=(\mathsf{depth}(C))^{2}\cdot\mathsf{qubit}(C)\enspace,$ (2)
@@
where $C$ is a circuit, $\mathsf{depth}(C)$ is the depth of $C$, and $\mathsf{qubit}(C)$ is the number of qubits required to run $C$. The function $S$ puts equal weights on time and space, whereas $T$ values time more than space as it requires quadratically more space to reduce time as in the quantum search algorithm [36]. Based on the two metrics $S$ and $T$, the cost of two extreme designs (the most space- and time-saving ones) for QFT-based additions is compared with the most cost-efficient circuit obtained by the algorithm. Table 1 summarizes the results. The circuit with the lowest cost is highlighted in Figure 3.
Table 1: Cost ratio of $C_{\rm space}$ and $C_{\rm time}$ to the most cost-efficient circuit obtained by the algorithm, where $C_{\rm space}$ and $C_{\rm time}$ are the most space- and time-saving circuits for QFT-based additions. Two different metrics $S$ and $T$ in Eq. (2) are examined. In Figure 3, in each tradeoffs curve, the circuits for $\min_{C}S(C)$ and $\min_{C}T(C)$ are highlighted by dashed circles and squares, respectively.
Figure 3 (a) Figure 3 (b)
$n=40$ $n=80$ $n=10$ $n=15$
$S(C_{\text{space}})/\min_{C}S(C)$ 1.21 1.33 1.33 1.48
$S(C_{\text{time}})/\min_{C}S(C)$ 2.53 3.50 3.85 4.16
$T(C_{\text{space}})/\min_{C}T(C)$ 3.03 5.14 6.90 12.27
$T(C_{\text{time}})/\min_{C}T(C)$ 1.73 2.19 1.93 1.82
5 Discussion
@@
In this work, optimizing the depth of a commuting circuit is reduced to solving graph coloring problem, leading to a strategy for circuit optimization. It has been shown that the strategy can be applied to improve the known circuits. Although this work concerns commuting circuits only, it is worth studying further applications as certain circuits can be transformed or divided into subcircuits appropriate for the proposed method. (Section 4.1 and Section 4.2 are such examples.)
@@
As a final note, we want to point out that the strategy developed in this work is capable of finding the optimal depth of a commuting circuit corresponding to a perfect graph [14], since there exists a polynomial time algorithm for graph coloring of a perfect graph [37]. The next question is then whether it is possible to find such a circuit that has practical uses while its optimal depth is not yet known. We have not formally proved yet, but considering that the perfect graphs do not have odd ring of length $\geq 5$, a circuit consisting of commuting CNOT gates seems to correspond to a perfect graph. [^3] On the other hand, CZ-only circuits generally do not correspond to perfect graphs. However, we were not able to take advantage of such a fact for a useful application.
Appendix 0.A Reduction from graph coloring to gate ordering
@@
$m_{\rm depth}({\bm{a}})=\min_{{\bm{a}}^{\prime}\in A_{\mathcal{V}}}\big{(}m_{\rm depth}({\bm{a}}^{\prime})\big{)}\enspace.$ (3)
@@
$m_{\rm depth}(f(b))\leq m_{\rm color}(b)<m_{\rm color}(g({\bm{a}}))=m_{\rm depth}({\bm{a}})\enspace,$
which contradicts to Eq. (3). $\square$

### v2

@@ 5.2 Depth-qubit tradeoffs for QFT-based addition
An algorithm for the parallelization by extra qubits can be devised by using graph coloring. A heuristic algorithm is developed in which the central idea is noted below. (We do not elaborate on the algorithm as the algorithm itself is not the key finding.)
@@
Consider a commuting circuit and its corresponding graph $G=(V,E)$. Since gates are vertices, introducing extra qubits leaves $V$ unchanged, but depending on which gates are lifted, $E$ changes such that the number of connected edges decreases. Figure 5 shows how the graph can be altered by one extra qubit. The key idea of the algorithm is that as we cut out the edges from the graph, the chromatic number can only get smaller. For example, let $G(n,p)$ be a random graph, where $n$ is the number of vertices and $p$ is the probability that an edge exists between two vertices. Bollobás has shown that the chromatic number of a random graph $\chi(G(n,p))$ asymptotically reads [56],
$\chi(G(n,p))\sim\frac{n}{2\log_{1/(1-p)}n}(1+o(1))\kern 5.0pt.$
In other words, the chromatic number of a random graph gets smaller as $p$ approaches 0. Therefore, we use the extra qubits in a way that the altered graph becomes as sparse in edge density as possible. [^3] Beware that not every commuting circuit corresponds to a random graph or a random intersection graph. By using the algorithm, we have parallelized the QFT-based additions $\lvert a\rangle\lvert b\rangle\mapsto\lvert a\rangle\lvert b+a\rangle$ and $\lvert a\rangle\lvert b\rangle\lvert c\rangle\mapsto\lvert a\rangle\lvert b\rangle\lvert c+ab\rangle$. For simplicity, we set $\lceil\log_{2}a\rceil=\lceil\log_{2}b\rceil=n$ and $c=0$. The results are summarized in Figure 6. (Up to roughly 10,000 qubits are considered, far exceeding what current hardware can support at the time of writing. One should take it as a theoretical calculation.)
Figure 6: Depth-qubit tradeoffs for QFT-based addition circuits (a) $\lvert a\rangle\lvert b\rangle\mapsto\lvert a\rangle\lvert b+a\rangle$ and (b) $\lvert a\rangle\lvert b\rangle\lvert c\rangle\mapsto\lvert a\rangle\lvert b\rangle\lvert c+ab\rangle$, where $a$ and $b$ are $n$-bit numbers and $c$ is set to zero. The circled and squared markers are the most efficient circuit designs with respect to the cost metrics in Eq. (3), respectively (see, Table 1).
Table 1: Cost ratio of $C_{\rm space}$ and $C_{\rm time}$ to the most cost-efficient circuit obtained by the algorithm, where $C_{\rm space}$ and $C_{\rm time}$ are the most space- and time-saving circuits for QFT-based additions. Two different metrics $S$ and $T$ in Eq. (3) are examined. The greater the value is than 1, the more inefficient it is. In Figure 6, in each tradeoffs curve, the circuits for $\min_{C}S(C)$ and $\min_{C}T(C)$ are highlighted by dashed circles and squares, respectively.
Figure 6 (a) Figure 6 (b)
$n=40$ $n=80$ $n=120$ $n=10$ $n=15$ $n=20$
$S(C_{\text{space}})/\min_{C}S(C)$ 1.21 1.33 1.41 1.33 1.48 1.57
$S(C_{\text{time}})/\min_{C}S(C)$ 2.53 3.50 5.02 3.85 4.16 5.57
$T(C_{\text{space}})/\min_{C}T(C)$ 3.03 5.14 6.66 6.90 12.27 18.84
$T(C_{\text{time}})/\min_{C}T(C)$ 1.73 2.19 2.77 1.93 1.82 2.12
@@
$S(C)=\mathsf{depth}(C)\cdot\mathsf{qubit}(C),\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ T(C)=(\mathsf{depth}(C))^{2}\cdot\mathsf{qubit}(C)\kern 5.0pt,$ (3)
@@
where $C$ is a circuit, $\mathsf{depth}(C)$ is the depth of $C$, and $\mathsf{qubit}(C)$ is the number of qubits required to run $C$. The function $S$ puts equal weights on time and space, whereas $T$ values time more than space as it requires quadratically more space to reduce time as in the quantum search algorithm [57]. Based on the two metrics $S$ and $T$, the cost of two extreme designs (the most space- and time-saving ones) for QFT-based additions is compared with the most cost-efficient circuit obtained by the algorithm. Table 1 summarizes the results. For example, in QFT-addition with $n=120$, $S(C_{\text{space}})/\min_{C}S(C)=1.41$ means the most space-saving circuit is 1.41 times more expensive than the optimized one found by the algorithm when considering the cost metric $S$. The circuit with the lowest cost is highlighted in Figure 6.
6 Conclusion and future directions
@@
In this work, optimizing the depth of a commuting circuit is reduced to solving graph coloring problem, leading to a strategy for circuit optimization. It has been shown that the strategy can be applied to improve the known circuits.
One limitation of the reduction strategy is that it only works on commuting circuits. Two workarounds can be further investigated. One direction is to transform a noncommuting circuit into one such that a portion of commuting parts is large. This is closely related with the circuit synthesis where the behavior of an input circuit should be maintained in the output [8, 26]. The other way is to divide a noncommuting circuit into many commuting parts and then applying the method where available. For example, LU decomposition is employed in optimizing linear circuits [58, 28], where we see a potential application of the coloring method.
@@
We want to point out that the strategy developed in this work is capable of finding the optimal depth of a commuting circuit corresponding to a perfect graph [14], since there exists a polynomial time algorithm for graph coloring of a perfect graph [59]. The next question is then whether it is possible to find such a circuit that has practical uses while its optimal depth is not yet known. We have not formally proved yet, but considering that the perfect graphs do not have odd ring of length $\geq 5$, a circuit consisting of commuting CNOT gates seems to correspond to a perfect graph. [^4] On the other hand, CZ-only circuits generally do not correspond to perfect graphs. However, we were not able to take advantage of such a fact for a useful application.
An anonymous reviewer raised a question whether the reduction idea can be generalized to account for hardware restrictions such as qubit connectivity and gate set. If hardware restrictions are to be considered, an originally commuting circuit will naturally become noncommuting one. We have tried to generalize the reduction idea so that it works on such restrictions, but came to the conclusion that it is highly nontrivial. There is a possibility that vertex coloring of a certain acyclic directed graph may properly reflect hardware restrictions, but we were not able to think of a clear and efficient way.
An anonymous reviewer raised another question if the vertex coloring approach is advantageous over the edge coloring of a hypergraph. In our estimation, it seems possible to find similar reduction steps by using the edge coloring in an appropriate hypergraph. To compare the better solvability of the edge coloring and the vertex coloring is an interesting but nontrivial task, since the target graphs as well as the known solvers are different. For example, the reduced hypergraph for edge coloring would have the same number of vertices and edges as the number of qubits and gates, respectively, whereas the reduced graph for vertex coloring has the same number of vertices as the number of gates but its edges are not directly relevant to the number of qubits. Resolving the above issues may be worth independent work.
Acknowledgement
We are grateful to Raghav Govind Jha for introducing the partitioning of Pauli strings and the relevant papers.
Appendix A Reduction from graph coloring to gate ordering
@@
$m_{\rm depth}({\bm{a}})=\min_{{\bm{a}}^{\prime}\in A_{\mathcal{V}}}\big(m_{\rm depth}({\bm{a}}^{\prime})\big)\kern 5.0pt.$ (4)
@@
$m_{\rm depth}(f(b))\leq m_{\rm color}(b)<m_{\rm color}(g({\bm{a}}))=m_{\rm depth}({\bm{a}})\kern 5.0pt,$
which contradicts to Eq. (4). ∎
Appendix B On remark 6
Case (a) in Remark 6 can only occurs when the coloring algorithm produces an approximate solution, not the true optimal one. Figure 8 (a) illustrates a commuting circuit and its corresponding graph. Having Assumption 9, the depth of the circuit is four. In Figure 8 (b), the graph is colored with the minimum number of colors, yielding the optimal solution. Consequently, the number of colors coincides with the depth of the resulting circuit. By contrast, Figure 8 (c) shows an approximate solution for the coloring. However, the depth of the resulting circuit is three since $v_{2}$ and $v_{4}$ happen to be simultaneously executable.
Figure 8: A commuting circuit and corresponding graph are shown in (a), and its colored graphs and resulting circuits are shown in (b) and (c).
Appendix C Coloring within Qiskit framework
Since the proposed method and Qiskit transpiler’s optimization are not directly comparable, we let the Qiskit compiler optimize commuting circuits, with and without the coloring method applied beforehand, into Clifford+$T$ circuits. By measuring the differences in two cases (with and without coloring applied), we assess how effectively the coloring enhances the compiler.
We randomly generate commuting circuits with $m=10$ qubits and $n\in\{10,20,...,100\}$ CCZ gates. Unlike the experiments in the main text, we fixed the size of the gates because when decomposing into a Clifford+$T$ circuit, imposing the same degree of parallelization on gates of different sizes (e.g., CCZ and CCCZ) is not well aligned with the goal of the experimental setup. The key Qiskit transpiler settings are as follows: OptimizationMetric is set to COUNT_T and optimization_level is set to 3. The results are shown in Figure 9.
Figure 9: Optimization by Qiskit transpiler with and without the coloring method applied for randomly generated commuting circuits with $m=10$ qubits and $n\in\{10,20,...,100\}$ gates. Markers with solid and dashed lines are output by Qiskit with optimization_level=0\;=0 and optimization_level=3\;=3, respectively.